# Bandit Learning in Housing Markets

### v1

@@ 6 Centralized Anytime Algorithm
While Algorithm 1 achieves a regret upper bound that matches the lower bound in Theorem 2 for all players, it requires prior knowledge of the time horizon $T$ to properly construct confidence intervals. Although one could apply the doubling trick (auer1995gambling) to make the algorithm anytime, this approach leads to non-monotonic expected instantaneous regret. Such behavior could raise concerns among players about the algorithm’s consistency during execution.
@@
Algorithm 2 Anytime Centralized Bandit Learning in Housing Markets
@@
$u^{(t)}(i,j)=\begin{cases}\infty&\text{if }T_{i,j}(t)=0,\\ \hat{{\bm{U}}}^{(t)}(i,j)+\sqrt{\frac{3\log t}{2T_{i,j}^{(t-1)}}}\quad&\text{otherwise.}\end{cases}$ (11)
Each player $p_{i}$ uses the index to compute a preference ranking $\hat{r}_{i,t+1}$, where arms are ordered by their UCBs in a decreasing order (e.g., $\mathop{\rm argmax}_{j}u^{(t)}(i,j)$ is ranked first).
@@
$Reg_{i}(T)\leq\sum_{j:\bar{\Delta}_{i,j}>0}\bar{\Delta}_{i,j}\left(\sum_{\mu\in M_{i,j}}\mathbb{E}T_{\mu}(T)\right).$ (12)
If a matching $\mu$ is not truly core-stable under the ground-truth utility matrix $\bm{U}$}, then a blocking coalition must exist. This coalition – a subset of players and their endowed arms – can reassign arms among themselves to improve every member’s utility. A blocking triplet $(p_{i},a_{j},a_{k})$ occurs when, under $\mu$, player $p_{i}$ is matched to $a_{j}$, yet the ground-truth utility satisfies ${\bm{U}}(i,k)>{\bm{U}}(i,j)$. Now, consider the set $M_{i,j}$ of all achievable matchings that are non-truly core-stable and where $p_{i}$ receives $a_{j}$. For this set, there exists a corresponding set of arms ${\mathcal{S}}_{i,j}\subseteq{\mathcal{A}}$ such that: for every $a_{k}\in\mathcal{S}{i,j}$, the triplet $(p_{i},a_{j},a_{k})$ is blocking for any $\mu\in M{i,j}$; while for any $a_{k^{\prime}}\in\mathcal{A}\setminus\mathcal{S}_{i,j}$, we have $\mathbf{U}(i,j)\geq\mathbf{U}(i,k^{\prime})$. Let $Q_{i,j}$ be the set of all such blocking triplets formed from ${\mathcal{S}}_{i,j}$. Crucially, for any implemented matching $\mu\in{\bf M}_{i,j}$, at least one triplet in $Q_{i,j}$ must be realized. This structure leads to the following regret upper bound.
@@
When all players rank arms according to the index defined in Eq.(11) and submit their preferences to the centralized platform, and the platform uses the TTC algorithm with the received preferences to compute an allocation each round, the core regret of player $p_{i}$ up to time $T$ satisfies
$Reg_{i}(T)$ $\leq\sum_{j:\bar{\Delta}_{i,j}>0}\bar{\Delta}_{i,j}\left[\sum_{(p_{i},a_{j},a_{k})\in Q_{i,j}}\left(5+\frac{6\log T}{\Delta_{i,k,j}^{2}}\right)\right]$
$\leq\max_{j}\bar{\Delta}_{i,j}\left(5N^{2}+12\frac{N\log T}{\Delta_{\min}^{2}}\right)$
$={\mathcal{O}}\left(\frac{N\log T}{\Delta_{\min}^{2}}\right)$
Theorem 3 offers a centralized problem-dependent ${\mathcal{O}}\left(\frac{N\log T}{\Delta_{\min}^{2}}\right)$ upper bound guarantees on the core regret of each player $p_{i}$, which also matches the lower bound derived in Section 5, and the proposed UCB-type algorithm is anytime and adaptive in the sense that the players do not need to know the time horizon $T$ and the minimum reward gap $\Delta_{\min}$ in advance.
@@
Let $T_{i,j,k}(T)$ be the number of times player $p_{i}$ pulls arm $a_{j}$ while the triplet $(p_{i},a_{j},a_{k})$ is blocking the matching selected by the platform. Since every time when a matching $\mu\in M_{i,j}$ is implemented, at least one of the blocking triplets in $Q_{i,j}$ happens, we have
$\sum_{\mu\in M_{i,j}}T_{\mu}(T)\leq\sum_{(p_{i},a_{j},a_{k})\in Q_{i,j}}T_{i,j,k}(T).$ (13)
By the definition of a blocking triplet, we know that if $p_{i}$ pulls arm $a_{j}$ when $(p_{i},a_{j},a_{k})$ is blocking, $p_{i}$ must have a higher upper confidence bound for $a_{j}$ than for $a_{k}$, i.e., we have the index $u(i,j)>u(i,k)$ while according to the ground-truth utility matrix we have ${\bm{U}}(i,j)<{\bm{U}}(i,k)$. Standard analysis for the single player UCB (bubeck2012regret) shows that
$\mathbb{E}T_{i,j,k}(T)\leq 5+\frac{6\log T}{\Delta^{2}_{i,k,j}}.$ (14)
@@
Furthermore, as there are at most $N$ arms such that $\bar{\Delta}_{i,j}>0$, we have
$\sum_{j:\bar{\Delta}_{i,j}>0}\bar{\Delta}_{i,j}\left[\sum_{(p_{i},a_{j},a_{k})\in Q_{i,j}}\left(5+\frac{6\log T}{\Delta_{i,k,j}^{2}}\right)\right]\leq\max_{j}\bar{\Delta}_{i,j}\left[\sum_{(p_{i},a_{j},a_{k})\in Q_{i,j}}\left(5N+\frac{6N\log T}{\Delta_{i,k,j}^{2}}\right)\right]$
When we consider the triplet set composed of all possible ${\bm{U}}(i,j)<{\bm{U}}(i,k)$, we have $\lvert Q_{i,j}\rvert\leq N$ and
$\sum_{k:{\bm{U}}(i,j)<{\bm{U}}(i,k)}\frac{1}{\Delta_{i,k,j}^{2}}\leq\sum_{k=1}^{N}\frac{1}{k^{2}\Delta_{\min}^{2}}\leq\frac{2}{\Delta_{\min}^{2}},$
which concludes the second inequality of the regret bound. ∎
@@ 7 Conclusion and Discussion
This paper introduces a novel online learning framework for housing markets with uncertain preferences, unifying two key objectives: core-stability and sample efficiency. We formalize this through core regret as our performance metric and propose two algorithms that bridge multi-armed bandit techniques with housing market mechanisms. These algorithms adapt to both centralized and decentralized settings, with guarantees for fixed-horizon and anytime environments. A matching regret lower bound proves the order-optimality of the algorithms, highlighting their theoretical and practical significance.
@@
This paper focuses on housing markets where each player initially possesses one endowed arm, maintaining an equal number of players and arms. However, real-world housing markets typically involve more complex scenarios: existing tenants with endowed arms coexist with new applicants lacking initial allocations, while vacant houses (free arms) may also be available (abdulkadirouglu1999house). In these generalized settings, the core losses its uniqueness, raising two fundamental challenges: first, identifying an appropriate tractable solution concept to serve as a benchmark, and second, determining whether efficient learning algorithms can achieve sublinear regret in this more realistic but complicated environment.
@@ Appendix B Proof of Regret Lower Bound
B.1 Divergence Decomposition
The proof of the divergence decomposition lemma builds upon the framework presented in Chapter 16 of lattimore2020bandit, extending it to accommodate multiple players. We first introduce some notations as follows.
@@ B.1 Divergence Decomposition
B.2 Proof of Regret Decomposition (Lemma 4)
@@
B.3 Proof of Regret Lower Bound (Theorem 2)
@@ Appendix C Technical Lemmas
Lemma 10 (Corollary 5.5 in lattimore2020bandit).
@@
Lemma 11 (Bretagnolle-Huber Inequality, Theorem 14.2 [lattimore2020bandit]).

### v2

@@ 6 Centralized Anytime Algorithm
While Algorithm 1 achieves a regret upper bound that matches the lower bound in Theorem 2 for all players, it requires prior knowledge of the time horizon $T$ to properly construct confidence intervals. Although one could apply the doubling trick (Auer et al. 1995) to make the algorithm anytime, this approach leads to non-monotonic expected instantaneous regret. Such behavior could raise concerns among players about the algorithm’s consistency during execution.
@@
Algorithm 2 Centralized Anytime UCB
@@
$u^{(t)}(i,j)=\begin{cases}\infty&\text{if }T_{i,j}(t)=0,\\ \hat{{\bm{U}}}^{(t)}(i,j)+\sqrt{\frac{3\log t}{2T_{i,j}(t-1)}},&\text{ otherwise.}\end{cases}$ (11)
Each player $p_{i}$ uses the index to compute a preference ranking $\hat{r}_{i,t+1}$, where arms are ordered by their UCBs in a decreasing order (e.g., $\operatorname*{arg\,max}_{j}u^{(t)}(i,j)$ is ranked first).
@@
$\begin{split}Reg_{i}(T)&\leq\sum_{j:\bar{\Delta}_{i,j}>0}\bar{\Delta}_{i,j}\left(\sum_{\mu\in M_{i,j}}\mathbb{E}T_{\mu}(T)\right)\\ &\leq\max_{j}\bar{\Delta}_{i,j}\left(\sum_{\mu\in M}\mathbb{E}T_{\mu}(T)\right),\end{split}$ (12)
where $M$ is the set of all achievable but non-truly core-stable matchings, i.e., $M=\cup_{(i,j)\in[N]\times[N]}M_{i,j}$. If a matching $\mu$ is not truly core-stable under the true utility matrix $\bm{U}$}, a blocking coalition must exist – a subset of players who can reassign their endowed arms to strictly improve every member’s utility. For any non-truly-core-stable matching $\mu\in M$, there exists a blocking coalition ${\mathcal{B}}_{\mu}$ such that players outside ${\mathcal{B}}_{\mu}$ keep their assignments under $\mu^{*}$. Within ${\mathcal{B}}_{\mu}$, at least one player $p_{m}$ must be part of a blocking triplet $(p_{m},a_{n},a_{n^{\prime}})$ where under $\mu$, $p_{m}$ is assigned $a_{n}$, ${\bm{U}}(m,n)<{\bm{U}}(m,n^{\prime})$ in truth yet the UCB index satisfies $u(m,n)>u(m,n^{\prime})$. If no such triplet existed, then for every matched pair $(p_{m},a_{n})$ and every $a_{n^{\prime}}$ with ${\bm{U}}(m,n)<{\bm{U}}(m,n^{\prime})$, we would have $u(m,n)<u(m,n^{\prime})$. Since $\mu$ is non-core-stable under $\bm{U}$}, there must exist a reallocation within ${\mathcal{B}}_{\mu}$ that strictly improves every member with respect to $\bm{U}$} – and, given the UCB ordering, also with respect to the UCB index. This would contradict the UCB-core-stability of the implemented matching. Let $Q_{m,n}$ be the collection of all such triplets, whenever a non-truly-core-stable $\mu\in M$ is implemented, at least one player-arm pair $(p_{m},a_{n})\in{\mathcal{B}}:=\cup_{\mu\in M}{\mathcal{B}}_{\mu}$ must realize a triplet from $Q_{m,n}$. This structure yields the following regret bound.
@@
Following Algorithm 2, the core regret of player $p_{i}$ up to time $T$ satisfies
$Reg_{i}(T)$ $\leq\max_{j}\bar{\Delta}_{i,j}\left[\sum_{\begin{subarray}{c}(p_{m},a_{n})\\ \in{\mathcal{B}}\end{subarray}}\sum_{\begin{subarray}{c}(p_{m},a_{n},a_{n^{\prime}})\\ \in Q_{m,n}\end{subarray}}\left(5+\frac{6\log T}{\Delta_{m,n^{\prime},n}^{2}}\right)\right]$
$\leq\max_{j}\bar{\Delta}_{i,j}\left(5N^{3}+12\frac{N^{2}\log T}{\Delta_{\min}^{2}}\right)={\mathcal{O}}\left(\frac{N^{2}\log T}{\Delta_{\min}^{2}}\right).$
Theorem 3 offers a centralized problem-dependent ${\mathcal{O}}\left(\frac{N^{2}\log T}{\Delta_{\min}^{2}}\right)$ upper bound guarantees on the core regret of each player $p_{i}$, and the proposed UCB-type algorithm is anytime and adaptive in the sense that the players do not need to know the time horizon $T$ and the minimum preference gap $\Delta_{\min}$ in advance.
@@
Let $T_{m,n,n^{\prime}}(T)$ be the number of times player $p_{m}$ pulls arm $a_{n}$ while $(p_{m},a_{n},a_{n^{\prime}})$ forms a blocking triplet. Since every time a matching $\mu\in M$ is implemented, at least one such blocking triplet $(p_{m},a_{n},a_{n^{\prime}})\in Q_{m,n}$ occurs for some $(p_{m},a_{n})\in{\mathcal{B}}$, we have
$\sum_{\mu\in M}T_{\mu}(T)\leq\sum_{\begin{subarray}{c}(p_{m},a_{n})\\ \in{\mathcal{B}}\end{subarray}}\sum_{\begin{subarray}{c}(p_{m},a_{n},a_{n^{\prime}})\\ \in Q_{m,n}\end{subarray}}T_{m,n,n^{\prime}}(T).$ (13)
From the analysis above, we know that when $(p_{m},a_{n},a_{n^{\prime}})$ is blocking, we have the UCB index $u(m,n)>u(m,n^{\prime})$ while ${\bm{U}}(m,n)<{\bm{U}}(m,n^{\prime})$ according to the ground-truth utility matrix. Standard analysis for the single player UCB (Bubeck et al. 2012) shows that
$\mathbb{E}T_{m,n,n^{\prime}}(T)\leq 5+\frac{6\log T}{\Delta^{2}_{m,n^{\prime},n}}.$ (14)
@@
When we consider the triplet set composed of all possible ${\bm{U}}(m,n)<{\bm{U}}(m,n^{\prime})$, we have $\lvert Q_{m,n}\rvert\leq N$ and
$\sum_{n^{\prime}:{\bm{U}}(m,n)<{\bm{U}}(m,n^{\prime})}\frac{1}{\Delta_{m,n^{\prime},n}^{2}}\leq\sum_{n^{\prime}=1}^{N}\frac{1}{(n^{\prime})^{2}\Delta_{\min}^{2}}\leq\frac{2}{\Delta_{\min}^{2}},$
and hence
$\max_{j}\bar{\Delta}_{i,j}\left[\sum_{(p_{m},a_{n})\in{\mathcal{B}}}\sum_{(p_{m},a_{n},a_{n^{\prime}})\in Q_{m,n}}\left(5+\frac{6\log T}{\Delta_{m,n^{\prime},n}^{2}}\right)\right]$
$\leq$ $\max_{j}\bar{\Delta}_{i,j}\left[\sum_{(p_{m},a_{n})\in{\mathcal{B}}}\left(5N+\frac{12\log T}{\Delta_{\min}^{2}}\right)\right].$
As there are at most $N^{2}$ possible player-arm pairs in the blocking coalition, the second inequality is concluded. ∎
@@ 7 Conclusion and Discussion
This paper introduces a novel online learning framework for housing markets with uncertain preferences, unifying two key objectives: core-stability and sample efficiency. We formalize this through core regret as our performance metric and propose two algorithms that bridge multi-armed bandit techniques with housing market mechanisms. These algorithms adapt to both centralized and decentralized settings, with guarantees for fixed-horizon and anytime environments. A matching regret lower bound proves the order-optimality of the decentralized algorithm.
@@
This paper focuses on housing markets where each player initially possesses one endowed arm, maintaining an equal number of players and arms. However, real-world housing markets typically involve more complex scenarios: existing tenants with endowed arms coexist with new applicants lacking initial allocations, while vacant houses (free arms) may also be available (Abdulkadiroğlu and Sönmez 1999). In these generalized settings, the core loses its uniqueness, raising two fundamental challenges: first, identifying an appropriate tractable solution concept to serve as a benchmark, and second, determining whether efficient learning algorithms can achieve sublinear regret in this more realistic but complicated environment.
@@ Appendix B Proof of Regret Lower Bound
Divergence Decomposition
The proof of the divergence decomposition lemma builds upon the framework presented in Chapter 16 of Lattimore and Szepesvári (2020), extending it to accommodate multiple players. We first introduce some notations as follows.
@@
Proof of Regret Decomposition (Lemma 4)
@@
Proof of Regret Lower Bound (Theorem 2)
@@ Appendix C Technical Lemmas
Lemma 10 (Corollary 5.5 in Lattimore and Szepesvári (2020)).
@@
Lemma 11 (Bretagnolle-Huber Inequality, Theorem 14.2 (Lattimore and Szepesvári 2020)).